A consumer agency randomly selected 1700 flights for two major airlines, A
and B. The following table gives the two-way classification of these flights
based on airline and arrival time. Note that ”less than 30 minutes late” in-
cludes flights that arrived early or on time.
Less Than 30 30 Minutes to More Than
Minutes Late 1 Hour Late 1 Hour Late
Airline A 429 390 92
Airline B 393 316 80
If one flight is selected at random from these 1700 flights, find the probability
that this flight is
(i) not more than 1 hour late

(ii) is not less than 30 minutes late

(iii) a flight on airline B given that it is 30 minutes to 1 hour late
1
Expert's answer
2021-09-09T23:57:06-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c: c}
 & Less \ than & 30\ minutes & More \ than & Total \\ 
 & 30 \ minutes& to\ 1\ hour & 1\ hour \\
 & late & late & 1\ hour \\
\hline
 Airlane\ A & 429 & 390 & 92 & 911 \\ 
 \hdashline
 Airlane\ B & 393 & 316 & 80 & 789 \\
 \hdashline
 Total & 822 & 706 & 172 & 1700 \\
\end{array}"


(i)


"P(\text{not more than 1hour late})=\dfrac{822+706}{1700}"

"=\dfrac{382}{425}"

(ii)


"P(\text{is not less than 30 minutes late})=1-\dfrac{822}{1700}"

"=\dfrac{439}{850}"

(iii)


"P(\text{ B | 30 minutes to 1hour late})=\dfrac{316}{706}"

"=\dfrac{158}{353}"